# Unsupervised Document and Template Clustering using Multimodal Embeddings

### v1

@@ Abstract
This paper investigates a novel approach to unsupervised document clustering by leveraging multimodal embeddings as input to traditional clustering algorithms such as $k$-Means and DBSCAN. Our method aims to achieve a finer-grained document understanding by not only grouping documents at the type level (e.g., invoices, purchase orders), but also distinguishing between different templates within the same document category. This is achieved by using embeddings that capture textual content, layout information, and visual features of documents. We evaluated the effectiveness of this approach using embeddings generated by several state-of-the-art pretrained multimodal models, including SBERT, LayoutLMv1, LayoutLMv3, DiT, Donut, and ColPali. Our findings demonstrate the potential of multimodal embeddings to significantly enhance document clustering, offering benefits for various applications in intelligent document processing, document layout analysis, and unsupervised document classification. This work provides valuable insight into the advantages and limitations of different multimodal models for this task and opens new avenues for future research to understand and organize document collections.
@@ 1 Introduction
Document clustering, the task of grouping documents into meaningful clusters without prior knowledge of their categories, plays a crucial role in organizing and understanding large document collections. Its applications span various domains, including information retrieval, document management, topic discovery, and, more recently, intelligent document processing (IDP). As the volume of digital documents continues to grow exponentially, efficient and accurate document clustering techniques have become increasingly important [1].
The field of document clustering has a rich history, and early approaches primarily relied on content-based features. Traditional methods often represented documents using the bag-of-words model, where each document is represented as a vector of word frequencies. The techniques like Term Frequency-Inverse Document Frequency (TF-IDF) were then used to weigh the importance of terms within the document collection, enabling the application of clustering algorithms such as $k$-Means and hierarchical clustering [2]. Although effective for broad topic-based clustering, these methods often fail to capture the semantic relationships between words and are sensitive to the vocabulary used.
The advent of deep learning has revolutionized many areas of natural language processing, including document clustering. Early deep learning approaches used autoencoders to learn lower-dimensional representations of documents, aiming to capture more semantic information than traditional methods [3]. More recently, the success of pretrained language models (LLM) such as BERT [4] has led to significant advancements in text-based document clustering. These models learn contextualized word embeddings, allowing for a richer understanding of document content. Techniques such as using embeddings from the [CLS] token or averaging word embeddings have become common practices for generating document-level representations for clustering [5]. Furthermore, recent work has explored the use of LLM embeddings directly for clustering, sometimes even leveraging the generative capabilities of LLMs to define cluster centroids [6, 7].
However, many types of documents possess inherent structural and visual characteristics beyond their textual content. For example, invoices, purchase orders, and reports, while belonging to different document categories, also exhibit distinct layouts and visual cues. Recognizing these patterns can be crucial for a more nuanced understanding and categorization of documents, particularly in the context of IDP and document layout analysis. Traditional text-based methods often overlook this valuable information, limiting their ability to distinguish between documents with similar textual content but different structures or templates.
@@
In this paper, we address this limitation by proposing a document clustering approach that uses multimodal embeddings. These embeddings are specifically designed to capture not only textual content but also layout structures and visual characteristics inherent in documents. By integrating these rich representations into standard clustering algorithms such as $k$-Means and DBSCAN, we aim to achieve a more comprehensive understanding of documents, facilitating clustering not only at the document-type level, but also at the template level within each type. For example, our approach can differentiate between distinct invoice templates from the same vendor or identify various versions of a financial report. Figure 1 illustrates the two granularity levels considered in our document clustering approach. At the document level (Figure 1(a)), documents are grouped according to their overall type, such as invoices, ID cards, and receipts. At a finer granularity, the template-level clustering (Figure 1(b)) separates documents of the same type based on variations in their layout or structure.
@@
To achieve this, we use embeddings from several state-of-the-art pretrained multimodal models, including SBERT (for text-only baseline) [8], LayoutLMv1 [9], LayoutLMv3 [10], DiT [11], Donut [12] and ColPali [13]. These models have demonstrated remarkable performance in various document understanding tasks by effectively integrating information from different modalities. By evaluating their embeddings for the task of unsupervised document clustering, we aim to provide insights into their strengths and weaknesses in capturing the nuances of document types and templates.
@@
  • We propose a novel approach for document clustering that leverages multimodal embeddings to group documents by both type and template.
  • We evaluated the effectiveness of embeddings from several state-of-the-art pretrained multimodal models (LayoutLMv1, LayoutLMv3, DiT, Donut, ColPali) for this task, along with a text-only baseline (SBERT).
@@
The remainder of this paper is structured as follows. Section 2 reviews prior research on deep learning-based document clustering, multi-view document clustering, and multimodal language models, situating our contribution within the existing literature. Section 3 details the proposed methodology: it explains how last-layer hidden states are projected into fixed-size embeddings, describes the six pretrained encoders we evaluate, and outlines the $k$-Means and DBSCAN algorithms used for clustering. Section 4 presents the empirical study: Section 4.1 formalizes the external and internal evaluation metrics, Sections 4.2 and 4.3 report template- and document-level clustering results on four heterogeneous corpora, and Section 4.4 analyzes the impact of pooling strategy and model scale. Finally, Section 5 synthesizes the findings.
@@ 2 Related Work
This section provides an overview of existing research related to our work. We categorize the related work into three main subsections: deep learning-based document clustering, multi-view document clustering, and multimodal language models.
2.1 Deep Learning-based Document Clustering
Deep learning has significantly impacted the field of document clustering, offering methods to learn more effective document representations. Diallo et al. [3] proposed a deep embedding clustering method based on a contractive autoencoder, with the objective of learning robust and discriminative embeddings for clustering. Another work explored representation learning for short text clustering, focusing on the challenges posed by the limited context in short documents [14]. Yan et al. [2] introduced a novel regularized concept factorization approach for document clustering, leveraging semantic concepts to improve clustering performance. In [15], a text document clustering method based on a weighted BERT model was presented, where the BERT attention mechanism is utilized to weight the importance of different words for clustering.
More recent work has focused on leveraging the power of large-language models (LLMs) for text clustering. The paper [6] investigates the use of embeddings from various LLMs for clustering different types of textual data. However, this study focuses mainly on text and does not evaluate the impact of incorporating layout and visual information through multimodal models. In contrast, our work specifically explores the benefits of multimodal embeddings for document clustering.
k-LLMmeans [7] proposes using LLM-generated summaries as cluster centroids, offering an interpretable and scalable approach to LLM-based text clustering. Although this method leverages the semantic understanding of LLMs, it remains focused on textual content. The foundational work on BERT [4] introduced a powerful pretrained language model that has been widely adopted for various NLP tasks, including generating document embeddings for clustering [5]. Xie et al. [16] proposed an unsupervised deep embedding method for clustering analysis, learning embeddings that are optimized for the clustering task itself.
A research work towards multi-modal, multi-task learning for document representation learning [17] explored pre-training a model on multiple modalities and tasks to learn better document representations. Although this work touches upon multimodality, our research specifically focuses on evaluating the effectiveness of existing pretrained multimodal models for the downstream task of unsupervised document clustering, particularly in distinguishing document templates.
2.2 Multi-View Document Clustering
Multi-view clustering aims to improve clustering performance by leveraging information from multiple distinct views or representations of the data. Several works have explored multi-view document clustering. Bai and Bai [18] proposed a method leveraging a multi-structure processor and a hybrid ensemble clustering module to improve multi-view document clustering. Zhu et al. [19] introduced multi-view deep subspace clustering networks to learn a common subspace for clustering data from different views. Brbić and Kopriva [20] explored multi-view low-rank sparse subspace clustering. Hu et al. [21] proposed a joint contrastive triple-learning approach for deep multi-view clustering. Bai et al. [22] presented a deep multi-view document clustering method with enhanced semantic embedding, and in another work, they proposed a hierarchical consensus learning model for deep multi-view document clustering [23]. A comprehensive survey on multi-view clustering is provided by Fang et al. [24]. Zhao et al. [25] introduced a double high-order correlation preserved robust multi-view ensemble clustering method. Yin et al. [26] proposed a shared generative latent representation learning approach for multi-view clustering.
While these works effectively utilize multiple views of documents, such as different textual features or representations derived from different algorithms, they often do not consider or fully exploit the views ensuing from multiple modalities like text, layout, and visual features in the same way that multimodal models do. Multimodal language models are specifically designed to learn joint representations by explicitly modeling the interactions and relationships between these different modalities. Therefore, leveraging the embeddings from these models for document clustering represents a distinct and relatively unexplored approach in the literature, which our research aims to investigate.
2.3 Multimodal Language Models
The field of multimodal language models has witnessed significant progress in recent years, with models capable of processing and understanding information from multiple modalities, including text, images, and layout. LayoutLMv1 [9] was one of the early models to effectively incorporate layout information into language understanding by jointly embedding text and layout. LayoutLMv2 [27] further enhanced this by adding visual features and introducing new pre-training tasks to better model the interactions between the three modalities. LayoutLMv3 [10] proposed a unified pre-training approach with masked language modeling and masked image modeling, achieving state-of-the-art results on various document AI tasks.
DiT (Document Image Transformer) [11] focused on self-supervised pre-training for document image understanding, demonstrating strong performance on vision-based document tasks. Donut (OCR-free Document Understanding Transformer) [12] presented an end-to-end model that directly processes document images without relying on explicit OCR, achieving competitive results. ColPali [13] introduced a vision-language model for efficient document retrieval, producing high-quality multi-vector embeddings from document images.
Other notable multimodal models include TILT (Text-Image-Layout Transformer) [28], which uses a unified transformer architecture to learn from text, image, and layout. TrOCR [29] focused on transformer-based optical character recognition. DocFormer [30] presented a multi-modal transformer architecture combining text, vision, and spatial features. While DocBERT [31] adapted BERT for document classification, it primarily focused on textual content. LiGT [32] explored layout-infused generative transformers for visual question answering on documents. Kosmos-2.5 [33] is a multimodal literate model excelling in understanding text-intensive images and generating structured markdown outputs. UDOP (Unifying Vision, Text, and Layout for Universal Document Processing) [34] introduced a foundation model for diverse document understanding and generation tasks. Unified-IO [35] presented a model capable of handling a wide range of AI tasks across vision, language, and multimodal domains. HRVDA (High-Resolution Visual Document Assistant) [36] addressed limitations in visual document understanding by efficiently processing high-resolution images.
While these multimodal models have shown great success in various document understanding tasks, their effectiveness for unsupervised document clustering, particularly in distinguishing different templates within the same document category, remains largely unexplored. Our work aims to bridge this gap by evaluating the clustering performance of embeddings from several of these state-of-the-art multimodal models. To the best of our knowledge, no previous work has specifically investigated the use of multimodal embeddings for clustering documents at both the type and template levels. This research therefore offers a novel perspective on leveraging the rich representations learned by these models for unsupervised document organization and understanding.
@@ 3.1 Document Representation via Last Hidden State Projections
Let $\mathbf{H}\in\mathbb{R}^{L\times D}$ denote the last hidden state tensor output from a transformer model for a single document, where $L$ is the sequence length and $D$ is the hidden dimension. To enable document-level clustering, we must reduce this 2D tensor to a single vector $\mathbf{v}\in\mathbb{R}^{d}$ (typically $d=D$ or smaller).
@@
This strategy preserves key textual and visual cues without excessively increasing dimensionality. Alternatively, dimensionality reduction can be achieved by principal component analysis (PCA) [37] on $\mathbf{H}_{t}$ and $\mathbf{H}_{v}$. Finally, an autoencoder can also be trained to reconstruct the concatenated representations $[\mathbf{H}_{t};\mathbf{H}_{v}]$, using its bottleneck representation as the document embedding. Although the latter approach delegates the optimal strategy for projecting the final hidden states to the trained model itself, it typically demands substantial amounts of training data and considerable computational resources.
@@ 3.2 Embedding Models
The selection of pretrained models is critical for effectively capturing the multimodal characteristics inherent in documents. Each model contributes uniquely by addressing distinct dimensions of document information. Sentence-BERT (SBERT) [8] serves as a text-only baseline; it represents a specialized variant of BERT designed explicitly to generate semantically meaningful sentence embeddings. Using SBERT allows one to evaluate the incremental benefit provided by incorporating layout and visual modalities into clustering tasks.
@@
The Document Image Transformer (DiT) [11], on the other hand, primarily targets visual aspects. DiT, a vision transformer pretrained through self-supervised learning on extensive document image datasets, allows the evaluation of the efficacy of visual representations independently of textual and layout contexts. Donut [12], an OCR-free Document Understanding Transformer, processes images directly without relying on external OCR modules. Using transformer-based architectures to generate textual content directly from visual inputs, it provides insights into the performance of end-to-end models.
Lastly, ColPali [13], a vision-language model tailored for efficient document retrieval, is trained to generate robust multi-vector embeddings that encapsulate both visual and textual semantics. Incorporating ColPali allows exploration of embeddings explicitly optimized for retrieval tasks within the clustering context.
Table 1: Overview of the models considered in this study, the input modalities exploited during pretraining, their parameter counts, and their encoder hidden sizes.
Model Modalities (pretraining) No. parameters Hidden size
SBERT (Base) Text 110M 768
@@
Table 1 summarizes the six encoders considered in this study, detailing their pretraining modalities, transformer hidden sizes, and parameter counts. Figure 3 illustrates how each pretrained model processes the document image during inference. Through this diverse and carefully curated selection of models, our objective is to thoroughly analyze how different modalities and architectures influence the effectiveness of document clustering according to document types and templates.
Figure 3: Document processing pipeline of each pretrained model during inference.
@@ 3.3 Clustering Algorithms
After obtaining the document embeddings from each of the selected models, we employ two widely used clustering algorithms: $k$-Means and DBSCAN.
@@ 3.3.1 $k$-Means
$k$-Means is a centroid-based partitioning algorithm that aims to divide $n$ samples into $k$ clusters by iteratively assigning each sample to the cluster with the nearest mean (centroid). The algorithm proceeds as follows:

### v3

@@ Abstract
We study unsupervised clustering of documents at both the category and template levels using frozen multimodal encoders and classical clustering algorithms. We systematize a model-agnostic pipeline that (i) projects heterogeneous last-layer states from text–layout–vision encoders into token-type–aware document vectors and (ii) performs clustering with centroid- or density-based methods, including an HDBSCAN + $k$-NN assignment to eliminate unlabeled points. We evaluate eight encoders (text-only, layout-aware, vision-only, and vision–language) with $k$-Means, DBSCAN, HDBSCAN + $k$-NN, and BIRCH on five corpora spanning clean synthetic invoices, their heavily degraded print-and-scan counterparts, scanned receipts, and real identity and certificate documents. The study reveals modality-specific failure modes and a robustness–accuracy trade-off, with vision features nearly solving template discovery on clean pages while text dominates under covariate shift, and fused encoders offering the best balance. We detail a reproducible, oracle-free tuning protocol and the curated evaluation settings to guide future work on unsupervised document organization.
@@ 1 Introduction
Document clustering groups documents into coherent sets without labels and underpins retrieval, management, topic discovery, and modern intelligent document processing (IDP). Its importance grows with the scale of digital collections [1]. Early work used content-only features—bag-of-words with Frequency-Inverse Document Frequency (TF–IDF) weighted vectors—clustered by $k$-means or hierarchical methods [2]. These approaches are effective for coarse topical structure but weak at capturing semantics.
Deep learning has advanced document representations. For example, autoencoders can learn compact embeddings [3]. Pre-trained language models such as BERT yield contextualized features, typically aggregated with the [CLS] token or with mean pooling to form document-level vectors [4, 5]. Building on this trend, recent studies use large language model (LLM) embeddings directly for clustering and even employ generation to propose cluster centroids [6, 7]. However, many documents carry a salient layout and visual signals beyond text. Invoices, purchase orders, and reports differ not only by category but also by template and page structure. Exploiting these cues is critical in IDP and document layout analysis, whereas text-only pipelines may conflate structurally distinct documents with similar wording.
@@
In this paper, we address this limitation by proposing a document clustering approach that uses multimodal embeddings. These embeddings are specifically designed to capture not only textual content but also layout structures and visual characteristics inherent in documents. By integrating these rich representations into standard clustering algorithms such as $k$-Means [8, 9, 10], DBSCAN [11], a combination of HDBSCAN [12] and $k$-NN [13], and BIRCH [14], we aim to achieve a more comprehensive understanding of documents, facilitating clustering not only at the document-type level, but also at the template level within each type. For example, our approach can differentiate between distinct invoice templates from the same vendor or identify various versions of a financial report. Figure 1 illustrates the two granularity levels considered in our document clustering approach. At the document level (Figure 1(a)), documents are grouped according to their overall type, such as invoices, ID cards, and receipts. At a finer granularity, the template-level clustering (Figure 1(b)) separates documents of the same type based on variations in their layout or structure.
@@
To achieve this, we use embeddings from several state-of-the-art pre-trained multimodal models, including SBERT (for text-only baseline) [15], LayoutLMv1 [16], LayoutLMv3 [17], DiT [18], Donut [19], ColPali [20], Gemma3 [21], and InternVL3 [22]. These models have demonstrated remarkable performance in various document understanding tasks by effectively integrating information from different modalities. By evaluating their embeddings for the task of unsupervised document clustering, we aim to provide insights into their strengths and weaknesses in capturing the nuances of document types and templates.
@@
  • A model-agnostic unsupervised pipeline for clustering documents at category and template granularity, including a token-type–aware projection for encoders that mix text and vision tokens, and a practical HDBSCAN + $k$-NN variant to obtain noise-free partitions without labels.
  • We evaluate the effectiveness of embeddings from several state-of-the-art pre-trained multimodal models (LayoutLMv1, LayoutLMv3, DiT, Donut, ColPali, Gemma3, and InternVL3) for this task, along with a text-only baseline (SBERT).
@@
The remainder of this paper is structured as follows. Section 2 reviews prior research on deep learning-based document clustering, multi-view document clustering, and multimodal language models, situating our contribution within the existing literature. Section 3 details the proposed methodology: it explains how last-layer hidden states are projected into fixed-size embeddings, describes the eight pre-trained encoders we evaluate, and outlines the algorithms used for clustering. Section 4 presents the empirical study. Section 4.1 describes the procedure adopted for hyperparameter selection. Section 4.2 formalizes the external and internal evaluation metrics. Sections 4.3 and 4.4 report template- and document-level clustering results on five heterogeneous corpora. Section 4.5 analyzes the impact of pooling strategy and model scale. Section 4.6 examines generalization to multi-page documents, situates our approach relative to weakly and semi-supervised alternatives, details post-clustering consolidation, and evaluates robustness to OCR noise and multilingual inputs. Finally, Section 5 synthesizes the findings.
@@ 2 Related Work
Deep Learning-Based Document Clustering.
Deep learning has advanced document clustering by learning representations rather than bag-of-words. Contractive autoencoders and deep embedding clustering improve robustness and clusterability [3, 23]. Additional lines of work address short-text representation learning [24], regularized concept factorization [2], and attention-weighted BERT for term importance [25]. Pre-trained language models such as BERT provide contextualized embeddings that are aggregated into document vectors [4, 5]. More recent studies employ LLM embeddings for clustering across domains [6] and use LLM-generated summaries as interpretable centroids [7]. Multimodal and multitask pretraining for document representations has also been explored [26]. In contrast, we evaluate off-the-shelf multimodal encoders for unsupervised document and template clustering, with a focus on layout and visual cues in addition to text.
Multi-View Document Clustering.
Multi-view clustering exploits complementary representations to improve grouping [27]. For documents, recent methods include a multi-structure processor with hybrid ensemble clustering [28], deep subspace approaches that learn shared latent spaces [29, 30], joint contrastive triple-learning [31], enhanced semantic embedding and hierarchical consensus learning for deep multi-view clustering [32, 33], robust ensemble modeling that preserves high-order correlations [34], and shared generative latent representations [35]. These techniques typically fuse alternative textual views or algorithmic representations rather than distinct modalities. In contrast, multimodal language models learn joint representations over text, layout, and visual signals. Using such multimodal embeddings for document clustering is a distinct and comparatively underexplored direction that we investigate in this work.
Multimodal Language Models.
Multimodal language models now jointly encode text, layout, and vision for document understanding. The LayoutLM family integrates layout in v1 and adds visual signals with expanded pretraining in v2, and v3 introduces a unified masked-language and masked-image objective [16, 36, 17]. Vision-focused models advance document image understanding and OCR-free processing, including DiT [18] and Donut [19]. ColPali produces multi-vector embeddings for retrieval [20]. General-purpose vision–language models such as Gemma3 [21] and InternVL3 [22] extend these capabilities with scalable pretraining and strong zero-shot performance. Additional architectures combine text, image, and spatial cues. TILT learns from all three modalities [37]. TrOCR targets transformer-based OCR [38]. DocFormer fuses text, vision, and spatial features [39]. DocBERT adapts BERT for document classification [40]. LiGT explores layout-infused generative reasoning [41]. Kosmos-2.5 parses text-intensive images [42]. UDOP and Unified-IO pursue broad multimodal processing [43, 44]. HRVDA scales to high-resolution documents [45]. Despite these advances, the suitability of multimodal embeddings for unsupervised document clustering, especially for discovering templates within a category, remains underexplored. We therefore evaluate the clustering performance of embeddings from several of these models for both document-type and template-level organization.
@@ 3.1 Document Representation via Last Hidden State Projections
Let $\mathbf{H}\in\mathbb{R}^{L\times D}$ denote the last hidden state tensor output from the encoder of a transformer model for a single document, where $L$ is the sequence length and $D$ is the hidden dimension. To enable document-level clustering, we must reduce this 2D tensor to a single vector $\mathbf{v}\in\mathbb{R}^{d}$ (typically $d=D$ or smaller).
@@
This strategy preserves key textual and visual cues without excessively increasing dimensionality. Alternatively, dimensionality reduction can be achieved by principal component analysis (PCA) [46] on $\mathbf{H}_{t}$ and $\mathbf{H}_{v}$. Finally, an autoencoder can also be trained to reconstruct the concatenated representations $\mathbf{H}_{t};\mathbf{H}_{v}]$, using its bottleneck representation as the document embedding. Although the latter approach delegates the optimal strategy for projecting the final hidden states to the trained model itself, it typically demands substantial amounts of training data and considerable computational resources.
@@ 3.2 Embedding Models
The selection of pre-trained models is critical for effectively capturing the multimodal characteristics inherent in documents. Each model contributes uniquely by addressing distinct dimensions of document information. Sentence-BERT (SBERT) [15] serves as a text-only baseline; it represents a specialized variant of BERT designed explicitly to generate semantically meaningful sentence embeddings. Using SBERT allows one to evaluate the incremental benefit provided by incorporating layout and visual modalities into clustering tasks.
@@
The Document Image Transformer (DiT) [18], on the other hand, primarily targets visual aspects. DiT, a vision transformer pre-trained through self-supervised learning on extensive document image datasets, allows the evaluation of the efficacy of visual representations independently of textual and layout contexts. Donut [19], an OCR-free Document Understanding Transformer, processes images directly without relying on external OCR modules. Using transformer-based architectures to generate textual content directly from visual inputs, it provides insights into the performance of end-to-end models.
ColPali [20], a vision-language model tailored for efficient document retrieval, is trained to generate robust multi-vector embeddings that encapsulate both visual and textual semantics. Incorporating ColPali allows exploration of embeddings explicitly optimized for retrieval tasks within the clustering context.
Gemma3 [21] is a multimodal vision–language model family that couples a language backbone with a SigLIP [47] visual encoder and a projection layer to align visual and language representations. In this work, we use the 4B variant: we take the image features after their projection into the language space and apply mean pooling to obtain a single embedding for clustering.
Among recent vision–language models, InternVL3 [22] represents a family that pairs a Qwen2.5 language backbone [48] with an InternViT visual encoder. In this work, we use the 14B variant, which couples Qwen2.5-14B with InternViT-300M. It is trained through a native multimodal pre-training paradigm, where the model is exposed to both text-only corpora and multimodal data simultaneously during one unified pre-training stage, rather than adapting a text-only LLM in a secondary phase. For clustering, we apply mean pooling over the projected image features derived from InternViT.
Table 1: Overview of the models considered in this study, the input modalities exploited during pre-training, their parameter counts, and their encoder hidden sizes.
Model Modalities (pre-training) No. parameters Hidden size
SBERT (Base) Text 110M 384
@@
Gemma3 Text, Image 4B 2560
InternVL3 Text, Image, Video 14B 1024
Table 1 summarizes the eight encoders considered in this study, detailing their pre-training modalities, transformer hidden sizes, and parameter counts. Figure 3 illustrates how each pre-trained model processes the document image during inference. Through this diverse and carefully curated selection of models, our objective is to thoroughly analyze how different modalities and architectures influence the effectiveness of document clustering according to document types and templates.
Figure 3: Document processing pipeline of each pre-trained model during inference.
@@ 3.3 Clustering Algorithms
After obtaining the document embeddings from each of the selected models, we employ widely used clustering algorithms as described next.
@@ 3.3.1 $k$-Means
$k$-Means [8, 9, 10] is a centroid-based partitioning algorithm that aims to divide $n$ samples into $k$ clusters by iteratively assigning each sample to the cluster with the nearest mean (centroid). The algorithm proceeds as follows: